OpenStudy (anonymous):

Find the equation for a parabola with vertex (-2, 0) and point on the graph (0, 1) [drawing below] I assume that the focus is (0, 0), so a = 2, so putting all these things into an equation I get: (y-0)^2 = 4(2)[x+2] y^2 = 8(x+2) But the book says it's y^2 = (1/2)(x+2) What am I doing wrong?

OpenStudy (anonymous):

Here's the graph shown in the book: 

Directrix (directrix):

The latus rectum is a segment so I'm wondering why you are giving it as if it were the point (0,1).

OpenStudy (anonymous):

Oh, sorry, I meant that it's a point on the graph in the same line as the focus. I don't the name for it, but I drew the graph from the book ^

OpenStudy (anonymous):

I've been working on it some more and I just realized the book never said that point is in line with the focus. But now when I try to solve for the focus, it sill doesn't work... y^2 = 4ax 1^2 = 4a(0) That gives me nothing!

Directrix (directrix):

But the book says it's y^2 = (1/2)(x+2) --------------------------------- General equation for horizontal parabola 4p ( x - h) = (y- k)² The vertex (h,k) is given to be (-2,0) Substitute that into the general formula. 4p ( x - (-2) ) = (y - 0)² 4p (x + 2) = y² (0,1) is a given point on the parabola Let (x1,y1) = (0,1) So, 4p ( (0) + 2 ) = ( (1) - 0) )² 4p(2) = 1 8p = 1 p = (1/8) ---- Returning to this earlier point: 4p (x + 2) = y² 4*(1/8) (x + 2) = y² (1/2) (x + 2) = y² y² = (1/2) (x + 2) @MGR Check this solution and see if you agree. Thanks.